OpenStudy (anonymous):

What is the solution to the system of equations? 2x+y=-3 x-2y=-4

OpenStudy (anonymous):

Re-arrange the 2nd equation: x - 2y = -4 Add 2y to both sides: x = 2y - 4 Plug in '2y - 4' for 'x' in the first equation: 2(2y - 4) + y = -3 Distribute 2: 4y - 8 + y = -3 5y - 8 = -3 Add 8 to both sides, what's -3 + 8?

OpenStudy (anonymous):

@becker2000

OpenStudy (anonymous):

idk im hurrying

OpenStudy (anonymous):

5